3x + 4 5x+6
_____ - ____
2x 3x

(3x+4 / 2x) - (5x+6 / 3x)

Answer by stanbon(75887)

3x + 4 5x+6
_____ - ____
2x 3x
======================================
lcd = 6x
-------------
(3x+4)/2x - (5x+6)/3x = (3)(3x+4) - (2)(5x+6)]/[6x]
-------------
[9x+12-10x-12]/6x
-------------
[-x/6x] = -(1/6)
=====================================
Cheers,
Stan H.
--------